### v1

# PainDiffusion: Can robot express pain?

Abstract
Pain is a more intuitive and user-friendly way of communicating problems, making it especially useful in rehabilitation nurse training robots. While most previous methods have focused on classifying or recognizing pain expressions, these approaches often result in unnatural, jiggling robot faces. We introduce PainDiffusion, a model that generates facial expressions in response to pain stimuli, with controllable pain expressiveness and emotion status. PainDiffusion leverages diffusion forcing to roll out predictions over arbitrary lengths using a conditioned temporal U-Net. It operates as a latent diffusion model within EMOCA’s facial expression latent space, ensuring a compact data representation and quick rendering time. For training data, we process the BioVid Heatpain Database, extracting expression codes and subject identity configurations. We also propose a novel set of metrics to evaluate pain expressions, focusing on expressiveness, diversity, and the appropriateness of model-generated outputs. Finally, we demonstrate that PainDiffusion outperforms the autoregressive method, both qualitatively and quantitatively. Code, videos, and further analysis are available at: https://damtien444.github.io/paindf/.
I INTRODUCTION
Pain is defined as ”an unpleasant sensory and emotional experience associated with actual or potential tissue damage.” The perception and expression of pain vary among individuals due to their unique life experiences [1]. In this work, we propose utilizing pain expressions in robots as a means of communicating system problems. Traditional error codes, such as a red number on a screen, can be obscure, easily misunderstood, and often require specialized knowledge to interpret. In contrast, pain expressions provide a more intuitive and universal method for indicating issues, enabling users of all backgrounds to understand and respond appropriately. Much like how animals, infants, and unconscious beings express distress, incorporating pain-like signals into robots creates a more user-friendly interface. Pain facial expressions can facilitate the immediate recognition of problems, which helps protect the system from further damage due to improper handling or adverse working conditions. By leveraging the inherent human understanding of pain, our approach aims to improve the interaction between users and robotic systems, ensuring timely and effective responses to potential issues.
Beyond this application, rehabilitation training robots must also express pain as feedback to nursing learners. During rehabilitation exercises, nurses must be able to read patients’ emotions to improve the quality of care. Previous works confirm that recognizing patients’ pain from their facial behavior enhances nursing skills and overall care quality [2, 3, 4]. Therefore, incorporating pain facial expressions in robotic patients can be valuable for nurse and doctor training, ultimately improving healthcare standards, especially for the subtle facial expressions that might help treatment more ease for patients.
In this work, our contribution is threefold. First, we propose PainDiffusion, a model designed to generate pain facial expressions that can handle arbitrary-length predictions, making it suitable for robotic applications. Second, we incorporate intrinsic characteristics, such as expressiveness and emotion, allowing for a more controllable generation tailored to diverse use cases. Third, we introduce a new set of baselines and metrics to effectively evaluate the quality and accuracy of pain expressions generated by our model.
Figure 1: Overview. PainDiffusion inputs pain stimuli signals, expressiveness configuration, emotion status, and past frames to generate the next appropriate prediction. By utilizing EMOCA latent as the facial expression representative, we effectively limit the data space, ensuring a smaller training time and inference time.
II RELATED WORKS
In rehabilitation nurse-training robots, earlier work primarily focused on recognizing pain situations and selecting expressions from a predefined set [4, 5]. However, this method often results in expressions that are unnatural and less effective. More recently, research has shifted towards using generative models for facial expressions [6, 7]. While non-likelihood generative models like Generative Adversarial Networks (GANs) have produced high-quality results across various domains, they face significant challenges due to the complexity and instability of training and optimization. Meanwhile, diffusion models and score-based models have emerged as powerful tools for generating images and videos. These models have now achieved state-of-the-art performance, particularly in generating realistic images and videos [8, 9, 10]. In human action generation, recent work is increasingly focusing on diffusion models to achieve state-of-the-art results [11, 12, 13].
Among pain-related research, most efforts have focused on recognition and classification using either traditional machine learning or deep learning techniques [14, 15, 16, 17]. PainDiffusion is the first method to explore pain expression generation using diffusion models, supporting arbitrary-length predictions.
III PAIN DIFFUSION
Our main goal is to model the relationship between facial expressions, pain-causing signals, and the intrinsic features of an individual, similar to the task of generating facial reactions based on a speaker’s speech or audio [6, 18]. To achieve this, we define the high-level task as follows: Given a continuous sequence of pain-causing signals and the configuration of the individual, we autoregressively predict the appropriate facial expression.
To capture ongoing reactions, we employ diffusion forcing (Sec. III-E) to roll a denoising diffusion model (Sec. III-D) to generalize the prediction further than the training temporal horizon. We developed a temporal latent U-Net (Sec. III-C) with temporal attention to enhance temporal coherence in the predictions. This model can process a sequence of conditioning signals and intrinsic configurations, generating a latent vector representing the output produced by EMOCA [19], as described in Sec. III-B. The high-level system is illustrated in Fig. 1.
III-A Problem Definition
Let $y\in\mathbb{R}^{d}$ represent a facial expression, where $Y=(y_{0},y_{1},\ldots,y_{n})$ is a sequence of facial expressions with length $n$. In our approach, the generation of $Y$ is conditioned on a pain-causing signal $C=(c_{0},c_{1},\ldots,c_{n})$, where each $c_{i}\in\mathbb{R}$. Additionally, the model is guided by a pain expressiveness configuration parameter $\mathcal{P}\in\mathbb{R}$, which controls the intensity of the pain expression, allowing the model to capture individual differences in how pain is expressed. Since a person’s pain expression may vary depending on their current emotional state, even under identical pain stimuli, we introduce an emotion configuration parameter $\mathcal{E}\in\mathbb{R}$. This parameter is included to adjust the generated facial expressions to account for the influence of the subject’s emotional state during the pain expression sequence.
III-B Facial Representations
We utilize EMOCA [19] to produce 3D latent codes as it effectively maps from a disentangled latent space to high-quality face meshes with a reasonable render time. EMOCA [19] builds on the FLAME [20] 3D face mesh model and DECA’s [21] method of decomposing an image $I$ into factors such as shape, albedo, and lighting, but places greater emphasis on maintaining emotion consistency in the output. The EMOCA latent representation is modeled as:
$E_{c}(I)\rightarrow(\boldsymbol{\beta},\boldsymbol{\theta},\boldsymbol{\psi},\boldsymbol{\alpha},\mathbf{l},\mathbf{c}),$ (1)
where $\boldsymbol{\beta}\in\mathbb{R}^{|\beta|}$ represents identity shape, $\boldsymbol{\theta}\in\mathbb{R}^{|\theta|}$ are pose parameters, $\boldsymbol{\psi}\in\mathbb{R}^{|\psi|}$ represents facial expressions, $\boldsymbol{\alpha}$ is albedo, $\mathbf{l}\in\mathbb{R}^{27}$ represents Spherical Harmonics lighting, and $\mathbf{c}\in\mathbb{R}^{3}$ represents camera parameters. In this approach, we focus specifically on facial expression and jaw pose, so the latent space is reduced to the concatenation $(\boldsymbol{\psi},\boldsymbol{\theta_{\text{jaw}}})$, while other features are fixed as the average values from the training dataset. By excluding less relevant information such as lighting conditions and other constant features, the model can concentrate on capturing the dynamics of facial expression changes.
After generating the new facial expression and jaw pose $(\boldsymbol{\psi},\boldsymbol{\theta_{\text{jaw}}})_{\text{pred}}$ , the 2D face image $I_{\text{pred}}$ can be rendered using the render function $R$ from PyTorch3D [22] along with the FLAME model $M$:
$\mathcal{R}(\boldsymbol{\psi},\boldsymbol{\theta_{\text{jaw}}})=R(M(\boldsymbol{\beta},\boldsymbol{\theta}_{\text{pred}},\boldsymbol{\psi}_{\text{pred}}),\boldsymbol{\alpha},\mathbf{l},\mathbf{c})\rightarrow I_{\text{pred}}.$ (2)
This rendering process allows the model to synthesize the predicted facial expression as a 2D image by leveraging the 3D mesh generated by the FLAME model. During the image render state, we learn that EMOCA latent space has a small variance for the pose parameter, which makes the diffusion model unable to focus on generating stable poses across frames, we scale the feature bigger to have the same variance with other parameters, then scale it down to render with EMOCA.
III-C Temporal Latent Unet
Latent diffusion models (LDMs) have been effectively utilized for generating images, videos, and human behaviors [23, 10, 13, 8, 24]. In our work, we follow the LDM approach by restricting data representation to the EMOCA latent space, resulting in a much smaller data space compared to conventional diffusion models for image or video generation. Prior works in video generation diffusion models [25, 23] have introduced temporal layers to the standard spatial U-Net [26] to better capture temporal information. Instead of employing 2D convolutions as in [10], we hypothesize that our facial latent space does not retain substantial structural information of $I$ after multiple layers of encoding in EMOCA [19]. Therefore, using 1D convolutions is sufficiently effective.
The network is designed with convolutional ResNet blocks [27], each followed by a spatial attention block and then a temporal attention block. The noise $\lambda_{t}$ is integrated into the ResNet block using scale and shift operators, while conditioning information $c=(C,\mathcal{P},\mathcal{E})$ is encoded with a lightweight MLP encoder and incorporated into the spatial attention block via cross-attention. To embed temporal information, we scale and shift the features using temporal embeddings. Both the temporal and noise-time embeddings are encoded using sinusoidal embeddings [28] for positional information. The up and down blocks have the architecture as illustrated in Fig. 2.
Figure 2: U-Net Blocks. The Temporal U-Net block incorporates $t_{\text{noise}}$ into the Convolution 1D ResNetBlock using a scale-shift operation. Next, spatial attention applies cross-attention to integrate the condition information $c$. Temporal time $t_{\text{temporal}}$ is embedded using sinusoidal embeddings, followed by cross-attention in the temporal attention block to help the model understand temporal dynamics. Finally, the features are scaled up or down. The skip connection in the up blocks is concatenated with $z$ from the down blocks.
Temporal Attention Layers. In general, this U-Net is similar to the standard spatial U-Net, with the key difference being the inclusion of temporal attention layers. Let $z\in\mathbb{R}^{B\times T\times C\times D}$ represent the video latent vector, where $D$ is the spatial latent dimension, $C$ is the channel, $T$ is time, and $B$ is the batch size. The spatial layers treat each frame independently as a batch of size $B\cdot T$, while the temporal layers operate on the temporal dimension, reinterpreting the latent vector as $z\in\mathbb{R}^{B\times C\times T\times D}$ for processing.
During training, we directly train the model using video data, rather than a mixture of images and videos as in [25, 23], because the latent space is compact enough to allow the model to learn both spatial and temporal features simultaneously. It is not necessary to predict every frame, especially those that are very close to each other, as they share most features. To optimize inference time, we adopt the frame stacking approach from [29] that pushes close frames to the channel dimension to generate simultaneously.
III-D Elucidated Diffusion
In this approach, we consider using the diffusion framework proposed by [30] as a more organized way to represent diffusion or score-based models, where the model is modeled as:
$D_{\theta}(\boldsymbol{z},\sigma)=c_{\text{skip}}(\sigma)\boldsymbol{z}+c_{\text{out}}(\sigma)\cdot F_{\theta}\bigg{(}c_{\text{in}}(\sigma)\boldsymbol{z},c_{\text{noise}}(\sigma),C\bigg{)},$ (3)
where $\boldsymbol{z}$ is the facial latent vector, $\sigma$ is the standard deviation of the Gaussian noise level, $C$ represents conditions, and $F_{\theta}$ is the temporal latent U-Net. By adjusting the terms $c_{\text{skip}},c_{\text{out}},c_{\text{in}},c_{\text{noise}}$, different diffusion strategies can be achieved with minimal changes to the model. We adopt the $c$ terms from the Elucidated Diffusion Model (EDM) [30] due to the flexibility of the network architecture. As a result, we do not consider reparameterization approaches, such as predicting velocity [31], since EDM [30] is a hybrid of both velocity, start, and noise prediction [32].
For clarity, the training objective is simplified as:
$\mathbb{E}_{\boldsymbol{y}\sim p_{\text{data}},\boldsymbol{n}\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}\mathbf{I}\right)}\bigg{[}\|\boldsymbol{y}_{\text{pred}}-\boldsymbol{y}_{\text{target}}\|_{2}^{2}\bigg{]},$ (4)
where
$y_{\text{pred}}=F_{\theta}\bigg{(}c_{\text{in }}(\sigma)\cdot(\boldsymbol{y}+\boldsymbol{n}),c_{\text{noise }}(\sigma)\bigg{)},$ (5)
and
$y_{\text{target}}=\frac{1}{c_{\text{out }}(\sigma)}\bigg{(}\boldsymbol{y}-c_{\text{skip }}(\sigma)\cdot(\boldsymbol{y}+\boldsymbol{n})\bigg{)}.$ (6)
To guide the generation and combine different condition signals, we add the following guidance [24, 33] during inference:
$\hat{z}=\left(1+\sum_{c\in C}\lambda_{c}\right)\cdot F_{\theta}\left(z,t,C\right)-\sum_{c\in C}\lambda_{c}\cdot F_{\theta}\left(z,t,\left.C\right|_{c=\emptyset}\right),$ (7)
where $\lambda_{c}$ is the guidance strength of the condition. During training, we randomly drop each condition with a probability of 0.1 and search for optimal guidance strength for each condition. This denoising diffusion model works on a short sequence video $Y$ because we consider physical pain facial expressions to be short-term behaviors, primarily focused on immediate stimuli signals and it speeds up the forward process. Therefore, we train the model on relatively short videos and conditions. We also randomly trim the conditions to account for the initial period when the condition is brief. During sampling, we use DPM++ [34] to support speed-up guided sampling that enables sampling high-quality samples within from 15 to 20 denoising steps.
III-E Diffusion Forcing
The goal of this model is to be applied to a robot and generate arbitrary-length predictions. We adopt diffusion forcing [29] to extend beyond the short training horizon. Diffusion forcing assigns different noise levels to each temporal timestep (Fig. 3), placing more uncertainty on future frames while reducing uncertainty for past frames. This allows the model to generate arbitrary-length predictions without divergence, as seen in [23]. Additionally, diffusion forcing eliminates the need for special guidance or extra forward passes to ensure temporal coherence, as required in [25]. It also introduces a hyperparameter that controls uncertainty—higher values result in greater uncertainty, which in turn requires more denoising steps.
Figure 3: Diffusion Forcing and Full Sequence Diffusion. Diffusion Forcing (a) applies different noise levels to each frame, whereas Full Sequence Diffusion (b) uses the same noise level for all frames.
As diffusion forcing requires, we train our model with random noise levels for each temporal frame. During sampling, we apply a scheduling matrix that denoises a window of frames $w$ and then shifts the window by a horizon step $h$, ensuring there are some overlapping context frames $w-h>0$. This overlap determines how quickly the model can respond to oncoming stimuli signals and how many past frames affect to the current generation. For example, with a small forward time, a sampling rate of 32Hz, and $h=16$, the model would react with a delay of 0.5 seconds and it consider 16 frames as context frames.
Figure 4: Qualitative comparison with Ground truth and Autoregressive baseline. Please visit the project website for better video qualitative results and more samples.
IV EXPERIMENTS
Our experiments are designed with two primary goals. First, we aim to demonstrate that our model outperforms the baseline and common approaches in facial expression generation. Second, we seek to show that our model can produce arbitrary-length predictions without divergence. To achieve these objectives, we propose a new set of metrics specifically for evaluating pain facial expression and long-term prediction errors.
IV-A Dataset
To create the pain facial expression dataset, we use the BioVid Heatpain Database part C [35, 36, 37], which captures heat pain responses from 87 subjects with 4 levels of pain intensity and 30 minutes of data, separated by pauses. The original dataset includes 3 modalities: frontal face video (25 FPS), biomedical signals, and heat stimuli signals. We split the dataset into two subsets: 61 subjects for training and 26 subjects for validation. The validation subsets maintain an equal ratio of male and female participants, and include 5 low-expression and 21 normal-expression subjects. We filter the original validation videos to contain sequences that are close to the moment the pain stimuli signal changes its intensity. We synchronize all the modalities to have the same sampling rate with the video.
Preprocessing. To compute the latent representation for each video frame in the dataset, we use EMOCA [19] to extract the expression code $\psi$ and jaw pose $\theta_{\text{jaw}}$, while calculating the mean face for all other features. To address the variance difference of $\theta_{\text{jaw}}$ being too small compared to $\psi$, which can cause jaw jiggling during generation, we scale $\theta_{\text{jaw}}$ up by 100 times to balance the loss term and ensure that both $\psi$ and $\theta$ are optimized more evenly. To determine the pain expression configuration for each subject, we calculate the Prkachin and Solomon Pain Intensity (PSPI) [38] by extracting action units (AUs) using the state-of-the-art GraphAU [39]. We then compute the average PSPI for each identity, referring to this as the pain expressiveness configuration. Additionally, we extract the emotion index for each subject using the HSEmotion emotion extractor [40], averaging across frames to create a consistent emotion configuration for each subject.
In its final form, the pain facial expression dataset consists of 4 modalities: facial latent vector, pain stimuli signal, pain expressiveness configuration, and emotion configuration. Our model is trained to generate the facial latent vector conditioned on the other modalities.
IV-B Experiments setup
Metrics. To compare the effectiveness of expressing pain, we draw from metrics used in multiple facial reaction generation [18, 6] and human behavior generation [13, 41, 42, 43] to propose the following set of evaluation metrics:
  • PainSim: Uses Dynamic Time Warping (DTW) to measure the temporal signal similarity between the generated sequence’s PSPI signal and the ground truth PSPI signal under the same pain stimuli.
  • PainCorr: Uses the Pearson Correlation Coefficient (PCC) to quantify the linear correlation between the generated PSPI signal and the ground truth PSPI signal.
  • PainAcc: Uses DTW to measure the similarity between the generated PSPI signal and the pain stimuli signal, verifying the relationship between the stimuli and the expression.
  • PainDist: Uses Pairwise Mean Squared Error (MSE) to evaluate the difference between the generated expressions and the ground truth expressions.
  • PainDivrs: MSE of multiple generated expressions under the same stimuli signals to assess the diversity of the generated outputs.
  • PainVar: Measures the variance of generated expressions within the same sequence to evaluate how varied the expressions are in a single sequence.
Baselines. We establish three baselines to validate the model’s effectiveness:
  • Nearest Neighbor: Performs segment search in the training dataset to find the pain stimuli signal most similar to the current signal.
  • Random Training Sequence: Returns a random sequence from the training dataset.
  • Vector Quantized VAE and Autoregressive Model: We test the effectiveness of the diffusion model compared to common approaches by applying the SOTA method in facial reaction generation from [44] on this dataset for comparison. For short, we refer to this method as autoregressive.
Implementation details. We train PainDiffusion with a sequence length of 64, a warm-up phase of 5k steps, and a total of 300k training steps. The model uses a frame stack of 4, 35 sampling steps, a learning rate of $4\times 10^{-4}$, and an exponential moving average (EMA) with a decay of 0.999. Training is conducted on a pair of NVIDIA 4060Ti and 3080 GPUs. All the metrics and qualitative output are computed with a generation length of 640 frames, 10 times longer than the training horizon, to confirm the model’s ability to generate arbitrary-length output.
IV-C Qualitative results
Fig. 4 shows a comparison of a sample with the same stimuli and the corresponding expressions generated by the autoregressive baseline, ground truth, and PainDiffusion. It demonstrates that the expressions generated by PainDiffusion are more natural and closer to the ground truth, while the autoregressive baseline tends to maintain a nearly static facial behavior. In the comparison between diffusion forcing and full-sequence diffusion, qualitatively, diffusion forcing produces more stable predictions, whereas full-sequence diffusion introduces greater diversity and randomness, with less of a lead-follow relationship with the stimuli signal. We provide non-cherry-picked samples on the project website for better qualitative representation.
To evaluate the magnitude and area of facial movements in the generated sequences, we use the Farneback method [45] to detect motion. By varying the stimuli, emotion configurations, and expressiveness configurations, the mean movements are shown in Fig. 5. Stronger stimuli levels result in stronger facial motion. Neutral emotion status shows the least facial movement, while sadness and happiness lead to larger movements when responding to pain stimuli. Higher pain expressiveness tends to result in weaker movements.
Figure 5: Average movement observed under varying stimuli levels, emotion configurations, and expressiveness configurations, while keeping other metrics constant. Higher stimuli levels correspond to greater movement, though different emotions exhibit varying levels of movement. Additionally, increased pain expressiveness tends to result in weaker overall movement.
Figure 6: PSPI [38] sequence of the Autoregressive baseline and PainDiffusion across a sequence and their movement area.
IV-D Quantitative results
TABLE I: Baselines. Comparison of our approach with the baseline.
PSPI [38] Expression Stimuli
PainSim ()($\downarrow$)( ↓ ) PainCorr ()($\uparrow$)( ↑ ) PainDist ()($\downarrow$)( ↓ ) PainDivrs ()($\uparrow$)( ↑ ) PainVar ()($\uparrow$)( ↑ ) PainAcc ()($\cdot$)( ⋅ )
Ground truth 0 0.9999 0.0000 0.0000 0.0552 17416
Random Training 342 0.5745 0.2273 0.1901 0.0163 17952
Nearest Neighbor 345 0.6773 0.2218 0.0000 0.0198 17864
FSQ-VAE Autoregressive [44] 297 0.3990 0.2338 0.0326 0.0159 17620
PainDiffusion w/ Full-seq Diffusion 216 0.4126 0.1578 0.0865 0.0454 18138
PainDiffusion w/ Diffusion Forcing 224 0.5839 0.1060 0.0638 0.0252 18110
()($\cdot$)( ⋅ ) means closer to ground truth is better.
Table I presents a comparison between our proposed method and other baseline approaches. Overall, our method outperforms the quantized-vector and autoregressive model across all metrics. Notably, our evaluation methods provide a more detailed perspective on the pain generation problem by using metrics based on the PSPI signal from both the generated output and the ground truth. The PSPI signal generated by PainDiffusion is more closely aligned with the ground truth compared to the heuristic and autoregressive baselines, as evidenced by lower PainSim and higher PainCorr.
In terms of expression diversity, PainDiffusion generates facial expressions that are not only closer to the ground truth but also exhibit greater variation across multiple predictions (lower PainDist, higher PainDivrs). Additionally, during a prediction sequence, our model achieves the highest expression variance, as indicated by higher PainVar. Despite this higher variance, analysis using the Farneback method [45] shows that PainDiffusion’s facial movements are more concentrated in key areas such as the mouth, eyebrows, chin, and nose (Fig. 6), which closely mirrors the ground truth (Fig. 4).
Regarding stimuli similarity, PainDiffusion achieves performance closer to the ground truth than the autoregressive baseline, with PainAcc being more aligned with the ground truth. As shown in Fig. 6, the stimuli signal tends to have a smoother shape, while the ground truth PSPI contains more fluctuations. It illustrates that PainDiffusion more closely follows the ground truth, particularly during periods of rapid stimuli change, indicating its sensitivity to dynamic pain variations. In contrast, autoregressive baseline shows less responsiveness to stimuli changes and deviates from the ground truth. Therefore, predictions that closely match the natural variability of the ground truth can be considered more accurate and realistic.
IV-E Ablation
TABLE II: Ablation Study Results. Comparison of different configurations.
PainSim PainCorr PainDist PainDivrs PainVar PainAcc
Context Window 2 322 -0.0110 0.1513 0.0771 0.0284 18302
Context Window 4 323 -0.0137 0.1337 0.0736 0.0278 18282
Context Window 8 332 -0.0059 0.1075 0.0678 0.0304 18286
DF Uncertainty 0.5 304 0.0195 0.1900 0.0847 0.0440 18296
DF Uncertainty 1 314 0.0265 0.1788 0.0795 0.0394 18299
DF Uncertainty 2 319 0.0332 0.1383 0.0743 0.0304 18290
DF Uncertainty 4 368 0.0224 0.1269 0.0684 0.0283 18325
Guiding 1_1_1 * 333 0.0147 0.1172 0.0634 0.0228 18285
Guiding 1_2_4 * 316 0.0033 0.1144 0.0651 0.0220 18273
Guiding 0.5_1_2 * 316 -0.0005 0.1267 0.0713 0.0276 18272
Guiding 0.25_0.5_1 * 329 0.0169 0.1354 0.0740 0.0278 18291
* Guiding strengths respectively are emotion configuration, pain expressiveness configuration, and pain stimuli signal.
We analyze the effect of different context sizes, diffusion forcing uncertainty, and guiding strength on performance, with the results presented in Tab. II. A longer context window results in a smaller expression distance from the ground truth (PainDist) and a better correlation with the PSPI signal (PainCorr). It also leads to greater expression variation across predictions. However, this comes with the tradeoff of increased prediction time, as a longer context window shortens the prediction horizon.
Regarding diffusion forcing uncertainty, this parameter controls the level of uncertainty the model considers when generating new windows. Higher uncertainty requires more steps for generation. Our results show that an uncertainty level of 2 produces the best correlation between the generated PSPI signal and the ground truth (PainCorr). Lower uncertainty levels yield better diversity (PainDivrs, PainVar) but result in greater deviation from the ground truth (PainDist).
In terms of guiding strength, we find that prioritizing stimuli signals, followed by pain expressiveness and then emotion configuration, produces the best results. Weaker guiding strength increases diversity (PainDivrs and PainVar) but sacrifices expression similarity to the ground truth (PainDist).
The choice of configuration can vary based on whether the focus is on the accuracy of pain expression levels or the diversity of pain expressions. Therefore, selecting the best configuration depends on which aspect—correctness or diversity—takes precedence in a given application.
V CONCLUSION
In this work, we presented PainDiffusion, a model designed to generate appropriate facial expressions in response to pain stimuli, with the ability to control pain expressiveness characteristics. PainDiffusion leverages diffusion forcing within a latent diffusion model that captures temporal information, enabling it to generate long-term predictions efficiently, making it suitable for robotic applications. The model generates more diverse and concentrated expressions compared to the autoregressive baseline, approaching the random baseline in terms of diversity while outperforming all baselines in terms of pain PSPI similarity and correlation.
The current approach focuses solely on non-verbal expressions and limits controllability to emotional status and pain expressiveness. Future work could explore additional parameters, such as wear characteristics of materials or actual physical damage, to generate more accurate pain expressions. Incorporating auditory pain expressions, such as sounds, could also provide a more natural and multimodal way of expressing pain.